# Implications of Identity of AI: Creators, Creations, and Consequences

### 29 Feb 2024 v1

@@ Abstract
The field of Artificial Intelligence (AI) is rapidly advancing, with significant potential to transform society. However, it faces a notable challenge: lack of diversity, a longstanding issue in STEM fields. In this context, This position paper examines the intersection of AI and identity as a pathway to understand biases, inequalities, and ethical considerations in AI development and deployment. We present a multifaceted definition of AI identity, which encompasses its creators, applications, and their broader impacts. Understanding AI’s identity involves analyzing the diverse individuals involved in AI’s development, the technologies produced, and the social, ethical, and psychological implications. After exploring the AI identity ecosystem and its societal dynamics, We propose a framework that highlights the need for diversity in AI across three dimensions: Creators, Creations, and Consequences through the lens of identity. This paper proposes the need for a comprehensive approach to fostering a more inclusive and responsible AI ecosystem through the lens of identity.
@@ Introduction
The concern is that while AI has immense potential for progress in areas like healthcare, agriculture, and education, its development is predominantly driven by the private sector (Abboud, Arya, and Pandi 2020), raising critical questions about its applicability and fairness (Dwivedi et al. 2021). Renowned AI safety groups and industry leaders have raised alarms not just about existential risks akin to pandemics and nuclear war but also about the social inequities AI might exacerbate (Hinton et al. 2023) (Buolamwini and Gebru 2018). Scholars have highlighted the immediate dangers of AI, especially its impact on marginalized and vulnerable populations (Birhane et al. 2022). This is evidenced by the shortcomings of existing AI models, including facial recognition inaccuracies and the limited effectiveness of Large Language Models (LLMs) in addressing the diverse needs of global populations (Coulter 2023). Congressional testimony on the need for AI regulation and the collective call by AI researchers for a pause in AI advancements emphasize the urgency of addressing these issues (Pause Giant 2023). Moreover, the public’s divided perception of AI, akin to debates in cybersecurity (Kovačević, Putnik, and Tošković 2020), points to a significant gap between the aspirational goals of various AI frameworks and principles, and the reality of their implementation, especially in diverse global contexts(Hagerty and Rubinov 2019).
@@
In this rapidly shifting landscape, a significant lack of diversity among its creators, researchers, and educators has emerged as a critical challenge that limits the range of innovative and inclusive AI solutions (Stathoulopoulos and Mateos-Garcia 2019). This evolving narrative and the need to solve the AI diversity gap underscores the need for a nuanced, socio-technical approach to understanding the multifaceted identity of AI (Lee and Qiufan 2021). Through an exploration of how AI is created, applied, and perceived (Kim et al. 2023) across various societal strata, as presented in the AI identity ecosystem (Figure 1), we can answer questions like, Who creates AI? Whose values does it reflect? How is it perceived across different cultures and demographics? And, how do all these factors influence the technology’s trajectory?(Bingley et al. 2023) It is intricately woven with the myriad societal dynamics that shape identity formation and influence perceptions. The answers to these questions are pivotal as they influence not only the technological trajectories of AI but also its societal implications as presented in the AI identity Framework (Figure 2). It is critical to underscore that the identity of AI isn’t simply a reflection of its perception but a pathway to bridging the AI divide and improving inclusivity efforts.
Figure 1: The AI Identity Ecosystem
@@
However, in understanding AI’s evolution it is essential to understand the confluence of various human endeavors that drive its development as mentioned above. In this section particularly we focus more on the identity of the creators as it has cascading effect on the other layers. These human contributors do more than just define AI’s trajectory; they embed societal norms, values, and biases into the technology, shaping the future of our automated world (Cave and Dihal 2020). The role of research cannot be overstated, as institutions constantly push the boundaries of what’s possible, often dictating the future direction of AI (Basole and Accenture 2021). Simultaneously, both tech behemoths and startups shape the AI landscape with their innovative products and platforms. Open-source communities bolster this growth, fostering collaboration and ensuring that AI tools remain accessible to all (Quan and Sanderson 2018). As the influence of AI permeates society, regulatory and ethical bodies are stepping in to ensure its deployment is responsible and aligned with societal values (Minkkinen, Zimmer, and Mäntymäki 2021). End-users and consumers, both individuals and businesses, are the final and important piece of the puzzle, adopting AI solutions and applications and influencing their development through feedback. Integral to all these components and often the ones that shape the direction of development is the financial pillar of investors and funding entities that fuel the growth of AI across sectors (Khanna 2023). Lastly, the AI ecosystem is enriched by educational institutions and platforms that impart knowledge, nurturing the current and future generations of AI enthusiasts and professionals through AI literacy (Long and Magerko 2020) and AI education (Schiff 2022).
@@
By examining the roles, interactions, and contributions of these diverse stakeholders as presented in the AI Ecosystem (Figure 1) we can gain a comprehensive understanding of how AI systems are created, operate, and evolve. By foregrounding these concepts, we aim to dissect how they influence AI creators and their creations. Through the depiction of layers and various pipelines (connections) across the layers that lead to various consequences or issues in the AI landscape, Figure 1 shows the myriad ways AI technologies can either reinforce existing societal disparities or potentially pave the way for a more inclusive and equitable future. By dissecting these relationships, we can better comprehend and address the disproportionate consequences and impacts that AI can have on various segments of society. The construct of human identity is influence as depicted in the image is present across the layers, present in the identities, data, perceptions, and experiences of end users in the AI ecosystem. Here Human Identity, as defined within intersectionality (Crenshaw 2013) and identity theory(Jenkins 2014), encompasses the multifaceted and interconnected aspects of an individual’s social categories, including race, gender (Scheuerman et al. 2020), class (Inaba and Togawa 2021), sexuality (Keyes, Hitzig, and Blell 2021), and disability (Trewin et al. 2019), nationality and age (Pollack 2005) etc, recognizing that these overlapping identities shape unique experiences of privilege and oppression (Benjamin 2019).
@@
Sociology Literature provides a comprehensive exploration of how individuals’ identities are formed (Stets and Burke 2000), negotiated (Jenkins 2014), and transformed (Collins and Bilge 2020) within the context of social structures and interactions, offering insightful analysis on the complexity of identity as both a personal and social construct (Albert 1998). This concept also extends to objects, organizations, and technology (Wood and Smith 2004), illustrating how identity formation transcends individual experiences to encompass broader societal interactions and constructs (Hirsch 1992). Historically, discussions and models involving identity in the AI landscape and ecosystem have been heavily centered on the technological and economic aspects (Mitchell 2020). With efforts focused on defining and explaining what AI is (Wang 2019) and how to understand its usage (Devedzic 2022), how AI interacts/perceives the human identity in various scenarios (Keyes, Hitzig, and Blell 2021)(Scheuerman et al. 2020). This includes discussions on algorithms (Noble 2018) (Pasquale 2015), data(Aragon et al. 2016) (Aragon et al. 2022), hardware infrastructure (Batra et al. 2019), application areas (Huang and Rust 2021), market growth, company roles, and investment trends (Mir, Kar, and Gupta 2022). While these aspects are undeniably crucial, they often overshadow the deeper, more intricate layers of AI’s relationship with society. We define ”AI Identity” in two dimensions: internal and external.
@@
  • Researchers, Scientists, and Academics: The development of AI is driven by a community of researchers, scientists, and academics from various disciplines (Hartmann and Henkel 2020). Experts in computer science, mathematics, biology, linguistics, cognitive science, and engineering collaborate to advance AI technologies. They conduct research, publish academic papers, and contribute to the theoretical foundations of AI(Stathoulopoulos and Mateos-Garcia 2019). Higher education institutions also play a big role in shaping the perception and standards (Mouta, Pinto-Llorente, and Torrecilla-Sánchez 2023)that these students carry into their professional careers. With HCI, interaction design, and UX communities contributing to AI work on addressing the gap in discourse about AI’s relationship to design practice  (Loi et al. 2019), higher education institutions and private research institutions are important actors in shaping the diversity in the field of AI (Ahmed and Wahed 2020).
@@
  • Collaborative Communities and Open-Source Projects: Collaboration and open-source initiatives play a crucial role in AI creation. Online communities, forums, and platforms facilitate knowledge sharing, collaboration, and the exchange of ideas among AI enthusiasts and professionals (Bostrom 2018) (Hibbard 2008). Open-source projects like TensorFlow and PyTorch provide accessible frameworks and tools that empower developers to create AI solutions (Paszke et al. 2019).
  • Government and Policy Institutions: Governments and policy institutions contribute to the creation of AI by establishing frameworks, regulations, and policies that guide its development and deployment (Bareis and Katzenbach 2022). The government also invests in the development of AI through its military/defense wing to be up to date with private system developments. Mulitaleral Global organizations in the ICT space are trying to address ethical considerations, and privacy concerns, and ensure responsible and accountable AI practices (Vacarelu 2022) (Floridi et al. 2021). Government-funded National research initiatives and grants also support AI research and innovation.
@@
To understand the definition of AI and its application in various scenarios it is important to understand the very building blocks of AI, and how they work and interact with the physical world (Raghavan 2021). This layer of the ecosystem acts as the mechanism or conduit that translate the goals reflected by the identity of the creators of AI into the consequences and considerations layer of AI Identity ecosystem, and therefore is key to understanding both the external and internal aspects of AI Identity. this should be explained in the beginning of this section and then summarized at the end of this section. In this paper, we will adopt a comprehensive approach to understanding the workings of artificial intelligence (AI) by analyzing it from both technical and socio-technical perspectives.
@@
  • Socioeconomic Impact: The implications of AI-driven automation have the potential to reshape job markets and socioeconomic structures through job displacement and disrupting the human resources landscape (Schellmann 2024). Thus, exploring strategies to ensure a just transition for workers and leveraging AI for socioeconomic development becomes crucial across various sectors like healthcare, education, and governance. Countries that rapidly adapt to AI stand to gain significant socioeconomic advantages, while those lagging behind face an exacerbated digital divide. This widening gap between AI-ready nations and those without adequate access or infrastructure could lead to significant disparities in social and economic growth in the coming decades escpecially for marginalised groups (Chetty et al. 2014).
@@
The perception of AI is not monolithic, It is influenced by numerous factors, including societal structures, personal experiences, media representations, etc. The lens of identity in the field adds a layer of complexity, which leads to discussions about whether AI application design and creation processes are informed by a myriad of backgrounds, experiences, and worldviews. Thus by understanding the context of who creates AI technologies, as well as examining the frameworks and societal considerations that guide AI’s development and the way they impact society, we can advocate for diversity and inclusion as essential to an AI Identity that serves all people fairly. The conceptual framework of AI identity ecosystem captures this complexity by examining the relationship between individual identity factors—such as race, ethnicity, gender, class, sexual orientation, religion, and disability—and their impact in technological contexts, specifically through the lens of creators and consumers roles in the development, access to, data management, perception, and implementation of AI in applications/solutions(Kivunja 2018). In the technology development landscape, the creator’s work is influenced by their experiences, perceptions and identity, which manifests in the data they select and collect, ultimately shaping the technology they create. Thus representation and inclusion in the creation process has far-reaching consequences and considerations, which the framework suggests must be critically assessed through the lens of identity. Key sociological constructs such as diversity, fairness, inclusion, and bias are interwoven with a fundamental sense of belonging, and accountability underscoring the importance of these concepts in evaluating diversity and inclusion work in the field of AI. The upward arrow alongside ’Impact of Identity’ suggests that the presence or lack of diversity and inclusion across the mentioned sociological constructs get amplified as we move up through the layers. for example, the impact of bias in the creator’s layers (Mehrabi et al. 2021) snowballs exponentially into the consequences and considerations layer much akin to the bio accumulation/magnification process in nature (Bommasani et al. 2022). Greater emphasis on these aspects can potentially elevate the role and positive influence of identity in the technological sphere. This sociological framework serves as a guide for a comprehensive analysis of how identity shapes technology and, conversely, how technology can reflect and affect societal values and individual sense of self.

### 10 Apr 2024 v2

# AI and Identity

@@ Abstract
AI-empowered technologies’ impact on the world is undeniable, reshaping industries, revolutionizing how humans interact with technology, transforming educational paradigms, and redefining social codes. However, this rapid growth is accompanied by two notable challenges: a lack of diversity within the AI field and a widening AI divide. In this context, This paper examines the intersection of AI and identity as a pathway to understand biases, inequalities, and ethical considerations in AI development and deployment. We present a multifaceted definition of AI identity, which encompasses its creators, applications, and their broader impacts. Understanding AI’s identity involves understanding the associations between the individuals involved in AI’s development, the technologies produced, and the social, ethical, and psychological implications. After exploring the AI identity ecosystem and its societal dynamics, We propose a framework that highlights the need for diversity in AI across three dimensions: Creators, Creations, and Consequences through the lens of identity. This paper proposes the need for a comprehensive approach to fostering a more inclusive and responsible AI ecosystem through the lens of identity.
@@ Introduction
The global population is currently around 8.1 billion. As of 2023, it’s estimated that there were roughly 3.4 billion people employed worldwide, with 2 billion people working within the informal economy (Dyvik 2023). The ICT sector was projected to employ 55.3 million people full-time by 2020, according to estimates made before the COVID-19 pandemic (Sherif 2023). Within this sector, a smaller number of workers are actively involved in developing or working with AI tools and systems. Looking forward, it’s expected that by 2025, up to 97 million individuals will be working in the AI field. Furthermore, the US AI market is anticipated to reach 299.64 billion dollars by 2026. AI tools and systems are projected to impact nearly 40 percent of jobs globally, with this figure rising to about 60 percent in advanced economies (Howarth 2024). These changes will lead to a combination of job replacement and augmentation with the influencing factor being access to or lack of access to the internet and digital infrastructure.
This scenario underscores a significant power imbalance, where a relatively small portion of the global population has a profound influence on the lives and livelihoods of the majority as visualized in Figure 1. In this rapidly evolving landscape and widening AI divide, a notable challenge that has emerged is the significant lack of diversity among the creators, researchers, and educators in the AI field. This homogeneity within the AI workforce, if it persists into the predicted 96 million AI workforce, represents more than an issue of fairness or representation; it fundamentally affects the design, implementation, and impact of AI technologies across different populations and societies. (Stathoulopoulos and Mateos-Garcia 2019).
@@
The concern is that while AI has immense potential for progress in areas like healthcare, agriculture, and education, its development is predominantly driven by the private sector (Abboud, Arya, and Pandi 2020) (Chui et al. 2023), raising critical questions about its applicability and fairness (Dwivedi et al. 2021). Renowned AI safety groups and industry leaders have raised alarms not just about existential risks akin to pandemics and nuclear war but also about the social inequities AI might exacerbate (Hinton et al. 2023) (Buolamwini and Gebru 2018). Scholars have highlighted the immediate dangers of AI, especially its impact on marginalized and vulnerable populations (Birhane et al. 2022). This is evidenced by the shortcomings of existing AI models, including facial recognition inaccuracies and the limited effectiveness of Large Language Models (LLMs) in addressing the diverse needs of global populations (Coulter 2023). Congressional testimony on the need for AI regulation and the collective call by AI researchers for a pause in AI advancements emphasize the urgency of addressing these issues (Pause Giant 2023). Moreover, the public’s divided perception of AI, akin to debates in cybersecurity (Kovačević, Putnik, and Tošković 2020), points to a significant gap between the aspirational goals of various AI frameworks and principles, and the reality of their implementation, especially in diverse global contexts(Hagerty and Rubinov 2019).
Figure 1: Representation of global workforce and AI industry impact
@@
This evolving narrative and the need to solve the AI diversity gap underscores the need for a nuanced, socio-technical approach to understanding the multifaceted identity of AI (Lee and Qiufan 2021). Through an exploration of how AI is created, applied, and perceived (Kim et al. 2023) across various societal strata, as presented in the AI identity ecosystem (Figure 2), we can answer questions like, Who creates AI? Whose values does it reflect? How is it perceived across different cultures and demographics? And, how do all these factors influence the technology’s trajectory?(Bingley et al. 2023) It is intricately woven with the myriad societal dynamics that shape identity formation and influence perceptions. The answers to these questions are pivotal as they influence not only the technological trajectories of AI but also its societal implications as presented in the AI identity Framework (Figure 2). It is critical to underscore that the identity of AI isn’t simply a reflection of its perception but a pathway to bridging the AI divide and improving inclusivity efforts.
Figure 2: The AI Identity Research Framework
@@
However, in understanding AI’s evolution it is essential to understand the confluence of various human endeavors that drive its development as mentioned above. In this section particularly we focus more on the identity of the creators as it has cascading effect on the other layers. These human contributors do more than just define AI’s trajectory; they embed societal norms, values, and biases into the technology, shaping the future of our automated world (Cave and Dihal 2020). The role of research cannot be overstated, as institutions constantly push the boundaries of what’s possible, often dictating the future direction of AI (Basole and Accenture 2021). Simultaneously, both tech behemoths and startups shape the AI landscape with their innovative products and platforms. Open-source communities bolster this growth, fostering collaboration and ensuring that AI tools remain accessible to all (Quan and Sanderson 2018).
As the influence of AI permeates society, regulatory and ethical bodies are stepping in to ensure its deployment is responsible and aligned with societal values (Minkkinen, Zimmer, and Mäntymäki 2021). End-users and consumers, both individuals and businesses, are the final and important piece of the puzzle, adopting AI solutions and applications and influencing their development through feedback. Integral to all these components and often the ones that shape the direction of development is the financial pillar of investors and funding entities that fuel the growth of AI across sectors (Khanna 2023). Lastly, the AI ecosystem is enriched by educational institutions and platforms that impart knowledge, nurturing the current and future generations of AI enthusiasts and professionals through AI literacy (Long and Magerko 2020) and AI education (Schiff 2022).
@@
By examining the roles, interactions, and contributions of these diverse stakeholders as presented in the AI Ecosystem (Figure 3) we can gain a comprehensive understanding of how AI systems are created, operate, and evolve. By foregrounding these concepts, we aim to dissect how they influence AI creators and their creations. Through the depiction of layers and various pipelines (connections) across the layers that lead to various consequences or issues in the AI landscape, Figures 1 and 2 together show the myriad ways AI technologies can either reinforce existing societal disparities or potentially pave the way for a more inclusive and equitable future.
By dissecting these relationships, we can better comprehend and address the disproportionate consequences and impacts that AI can have on various segments of society. The construct of human identity is influenced as depicted in the image is present across the layers, present in the identities, data, perceptions, and experiences of end users in the AI ecosystem. Here Human Identity, as defined within intersectionality (Crenshaw 2013) and identity theory(Jenkins 2014), encompasses the multifaceted and interconnected aspects of an individual’s social categories, including race, gender (Scheuerman et al. 2020), class (Inaba and Togawa 2021), sexuality (Keyes, Hitzig, and Blell 2021), and disability (Trewin et al. 2019), nationality and age (Pollack 2005) etc, recognizing that these overlapping identities shape unique experiences of privilege and oppression (Benjamin 2019).
@@
Sociology Literature provides a comprehensive exploration of how individuals’ identities are formed (Stets and Burke 2000), negotiated (Jenkins 2014), and transformed (Collins and Bilge 2020) within the context of social structures and interactions, offering insightful analysis on the complexity of identity as both a personal and social construct (Albert 1998). This concept also extends to objects, organizations, and technology (Wood and Smith 2004), illustrating how identity formation transcends individual experiences to encompass broader societal interactions and constructs (Hirsch 1992). Often discussions and models involving identity in the AI landscape and ecosystem are heavily centered on the technological and economic aspects (Mitchell 2020). With efforts focused on defining and explaining what AI is (Wang 2019) and how to understand its usage (Devedzic 2022) (Touretzky, Gardner-McCune, and Seehorn 2023), how AI perceives the human identity (Scheuerman et al. 2020) (Schlesinger, O’Hara, and Taylor 2018) (Tian et al. 2017), how humans perceive AI (Ragot, Martin, and Cojean 2020) (Lima et al. 2020) (Shinners et al. 2022), how humans interact with the AI (Keyes, Hitzig, and Blell 2021) (Ashktorab et al. 2020) and how AI influences the human identity in various scenarios (Cao et al. 2023). This includes discussions on algorithms (Noble 2018) (Pasquale 2015), data(Aragon et al. 2016) (Aragon et al. 2022), hardware infrastructure (Batra et al. 2019), application areas (Huang and Rust 2021), market growth, company roles (Alahmad and Robert 2020), and investment trends (Mir, Kar, and Gupta 2022). While these aspects are undeniably crucial, they often overshadow the deeper, more intricate layers of AI’s perception and relationship with the concept of identity.
We define ”AI Identity” in two dimensions: internal and external.
@@
  • Researchers, Scientists, and Academics: The development of AI is driven by a community of researchers, scientists, and academics from various disciplines (Hartmann and Henkel 2020). Experts in computer science, mathematics, biology, linguistics, cognitive science, and engineering collaborate to advance AI technologies. They conduct research, publish academic papers, and contribute to the theoretical foundations of AI(Stathoulopoulos and Mateos-Garcia 2019). Higher education institutions also play a big role in shaping the perception and standards (Mouta, Pinto-Llorente, and Torrecilla-Sánchez 2023)that these students carry into their professional careers. With HCI, interaction design, and UX communities contributing to AI work on addressing the gap in discourse about AI’s relationship to design practice  (Loi et al. 2019), higher education institutions and private research institutions are important actors in shaping the diversity in the field of AI (Ahmed and Wahed 2020) (Whittaker et al. 2018).
@@
  • Collaborative Communities and Open-Source Projects: Collaboration and open-source initiatives play a crucial role in AI creation. Online communities, forums, and platforms facilitate knowledge sharing, collaboration, and the exchange of ideas among AI enthusiasts and professionals (Bostrom 2018) (Hibbard 2008). Open-source projects like TensorFlow and PyTorch provide accessible frameworks and tools that empower developers to create AI solutions (Paszke et al. 2019). Non-profits and professional societies are also getting involved with building standards and expectations around AI integration from the lenses of ethics and standards.
  • Government and Policy Institutions: Governments and policy institutions contribute to the creation of AI by establishing frameworks, regulations, and policies that guide its development and deployment (Bareis and Katzenbach 2022) (Luo and Van Assche 2023). The government also invests in the development of AI through its military/defense wing to be up to date with private system developments. Mulitaleral Global organizations in the ICT space are trying to address ethical considerations, and privacy concerns, and ensure responsible and accountable AI practices (Vacarelu 2022) (Floridi et al. 2021). Government-funded National research initiatives and grants also support AI research and innovation (Donlon 2024).
@@
To understand the definition of AI and its application in various scenarios it is important to understand the very building blocks of AI, and how they work and interact with the physical world (Raghavan 2021). This layer of the ecosystem acts as the mechanism or conduit that translate the goals reflected by the identity of the creators of AI into the consequences and considerations layer of AI Identity ecosystem, and therefore is key to understanding both the external and internal aspects of AI Identity. this should be explained in the beginning of this section and then summarized at the end of this section. In this paper, we will adopt a comprehensive approach to understanding the workings of artificial intelligence (AI) by analyzing it from both technical and socio-technical perspectives. It is simultaneously important to understand the intricate connection between the infrastructure layer and the growth of AI’s impact on society through these two lenses.
@@
This infrastructure layer is not only a precursor to the technical layer’s growth but also a fundamental aspect that interweaves with the socio-technical layer, affecting how AI integrates and resonates within human society. Through its capacity to support and scale AI development, It not only provides the technical means for AI’s existence but also influences the objectives, ethics, and inclusivity of AI systems by virtue of access. By enabling or limiting the development of certain AI applications, the infrastructure layer shapes the AI identity, reflecting the priorities and values of its creators and users. Furthermore, by offering metrics for capacity building, the infrastructure layer facilitates the assessment of global AI development efforts, guiding policy, investment, and education towards a more equitable and inclusive AI future.
The Infrastructure Layer
The infrastructure underpinning artificial intelligence (AI) is the bedrock upon which the technical evolution and societal integration of AI is built. This infrastructure encompasses the essential tools, hardware, and frameworks necessary for the development, training, and deployment of AI systems, including data processing capabilities, computational power, and cloud services. These components are indispensable for the processing of vast datasets, supporting complex model training, and ensuring the scalable deployment of AI applications. As such, the infrastructure directly influences the pace and direction of AI innovation. The disparities in AI capabilities across different regions highlight the need for global cooperation and investment in AI infrastructure. Consequently, infrastructure availability becomes a crucial metric for gauging capacity-building efforts, underscoring the importance of democratizing AI benefits to mitigate regional disparities. As AI technologies become increasingly embedded in everyday life, the infrastructure that supports these technologies plays a vital role in determining how they are received, understood, and integrated into various cultural contexts. This interaction between AI and society influences the design, functionality, and adoption of AI systems, necessitating a thoughtful approach to infrastructure development that considers and reflects diverse human values and societal needs.
  • Large Language Models (LLMs): Large Language Models, such as GPT (Generative Pretrained Transformer) and BERT (Bidirectional Encoder Representations from Transformers), represent significant advancements in natural language processing and understanding. These models, which learn from vast amounts of text data, have revolutionized how machines understand and generate human-like text, enabling a wide range of applications from automated writing assistants to advanced chatbots. The development and refinement of LLMs require substantial computational resources and sophisticated algorithms, showcasing the importance of robust infrastructure in AI innovation (Devlin et al. 2018) (Brown et al. 2020).
  • Data: Data serves as the cornerstone for training AI models, providing the raw material from which machines learn and make inferences. The quality, diversity, and volume of data directly impact the performance and bias in AI systems. Effective data management and processing infrastructure are crucial for handling the increasing scale of data, necessitating advanced storage solutions and data processing frameworks (Hartmann and Henkel 2020). The ethical considerations in data collection and use are also paramount, emphasizing the need for transparency and fairness in AI (Aragon et al. 2022).
  • Curriculum: The AI curriculum encompasses the educational resources and programs designed to equip individuals with the knowledge and skills required to develop and manage AI technologies. This includes courses on machine learning, ethics in AI, data science, and specialized AI applications. The development of a comprehensive AI curriculum that is accessible and inclusive is essential for fostering a diverse and skilled workforce capable of advancing AI technology while considering its societal impacts (Maher and Tadimalla 2024) (Long and Magerko 2020) (Schiff 2022) (Song et al. 2024).
  • Standards and Policy: Standards and policies play a critical role in guiding the ethical development, deployment, and governance of AI technologies. These frameworks help ensure that AI systems are developed responsibly, promoting transparency, accountability, and fairness. International organizations and regulatory bodies are increasingly focusing on developing AI standards and policies that balance innovation with ethical considerations and human rights (Jobin, Ienca, and Vayena 2019) (Long et al. 2023)
  • Hardware: The hardware infrastructure, including GPUs (Graphics Processing Units) and TPUs (Tensor Processing Units), is fundamental for training and running AI models (Jouppi et al. 2017). These specialized processors can handle the parallel processing of large datasets and complex algorithms, making them indispensable for deep learning applications. The development of more efficient and powerful hardware is crucial for advancing AI capabilities and making AI more accessible and sustainable (Merritt 2024).
  • Cloud and Digital Tools: Cloud computing and digital tools offer scalable and flexible resources for AI development and deployment. Cloud services provide access to computational power, storage, and AI development tools on-demand, facilitating the rapid prototyping and scaling of AI applications. This infrastructure supports the collaborative development of AI, enabling access to cutting-edge technology without significant upfront investment (Armbrust et al. 2010).
In essence, the infrastructure supporting AI is not just a foundation for its technical growth but also a mediator of its interaction with society, influencing both the capabilities of AI technologies and their alignment with human values.
@@
  • Socioeconomic Impact: The implications of AI-driven automation have the potential to reshape job markets and socioeconomic structures through job displacement and disrupting the human resources landscape (Schellmann 2024). Thus, exploring strategies to ensure a just transition for workers and leveraging AI for socioeconomic development becomes crucial across various sectors like healthcare, education, and governance (Chui et al. 2023). Countries that rapidly adapt to AI stand to gain significant socioeconomic advantages, while those lagging behind face an exacerbated digital divide. This widening gap between AI-ready nations and those without adequate access or infrastructure could lead to significant disparities in social and economic growth in the coming decades escpecially for marginalised groups (Chetty et al. 2014).
@@
The perception of AI is not monolithic, It is influenced by numerous factors, including societal structures, personal experiences, media representations, etc. The lens of identity in the field adds a layer of complexity, which leads to discussions about whether AI application design and creation processes are informed by a myriad of backgrounds, experiences, and worldviews. Thus by understanding the context of who creates AI technologies, as well as examining the frameworks and societal considerations that guide AI’s development and the way they impact society, we can advocate for diversity and inclusion as essential to an AI Identity that serves all people fairly. The research and conceptual frameworks of AI identity ecosystem capture this complexity by examining the relationship between individual identity factors—such as race, ethnicity, gender, class, sexual orientation, religion, and disability—and their impact in technological contexts, specifically through the lens of creators and consumers roles in the development, access to, data management, perception, and implementation of AI in applications/solutions(Kivunja 2018). In the technology development landscape, the creator’s work is influenced by their experiences, perceptions and identity, which manifests in the data they select and collect, ultimately shaping the technology they create. Thus representation and inclusion in the creation process has far-reaching consequences and considerations, which the framework suggests must be critically assessed through the lens of identity. Key sociological constructs such as diversity, fairness, inclusion, and bias are interwoven with a fundamental sense of belonging, and accountability underscoring the importance of these concepts in evaluating diversity and inclusion work in the field of AI. The upward arrow alongside ’Impact of Identity’ suggests that the presence or lack of diversity and inclusion across the mentioned sociological constructs get amplified as we move up through the layers. for example, the impact of bias in the creator’s layers (Mehrabi et al. 2021) snowballs exponentially into the consequences and considerations layer much akin to the bio accumulation/magnification process in nature (Bommasani et al. 2022). Greater emphasis on these aspects can potentially elevate the role and positive influence of identity in the technological sphere. This sociological framework serves as a guide for a comprehensive analysis of how identity shapes technology and, conversely, how technology can reflect and affect societal values and individual sense of self.